OpenStudy (anonymous):

A toy car in a science experiment covers 1.6 meters in half a second. If the car travels at a steady speed, how far will it go in 10 seconds? A. 0.8 m B. 8 m C. 16 m D. 32 m

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@johnweldon1993 @Chelsey2015 @coco_blue @GeniousCreation

OpenStudy (johnweldon1993):

Hint...if it covers 1.6m in (1/2) seconds...then in 10 seconds at the same speed it will cover (x) It's a proportion \[\large \frac{1.6m}{0.5s} = \frac{x}{10s}\] and solve for 'x'

OpenStudy (anonymous):

@johnweldon1993 so the answer would be 16

OpenStudy (johnweldon1993):

Not quite...when we cross multiply we get [\[\large 10\times 1.6 = 0.5x\] \[\large 16 = 0.5x\] What would 'x' be?

OpenStudy (anonymous):

so the answer would be 8

OpenStudy (johnweldon1993):

Not right again :/ \[\large 16 = 0.5x\] ^ "One half of SOME number = 16*

OpenStudy (anonymous):

Oh 32

OpenStudy (johnweldon1993):

There we go :)